# Experimental verification of the quantum nature of a neural network

### v3

@@ Abstract
Neural networks are being used to improve the probing of the state spaces of many particle systems as approximations to wavefunctions and in order to avoid the recurring sign problem of quantum monte-carlo. One may ask whether the usual classical neural networks have some actual hidden quantum properties that make them such suitable tools for a highly coupled quantum problem. I discuss here what makes a system quantum and to what extent we can interpret a neural network as having quantum remnants. I suggest that a system can be quantum both due to its fundamental quantum constituents and due to the rules of its functioning, therefore, we can obtain entanglement both due to the quantum constituents’ nature and due to the functioning rules, or, in category theory terms, both due to the quantum nature of the objects of a category and of the maps. From a practical point of view, I suggest a possible experiment that could extract entanglement from the quantum functioning rules (maps) of an otherwise classical (from the point of view of the constituents) neural network.
@@ I Introduction
Probably the best way of starting this article is to ask a question: what means to be quantum? Lacking a proper axiomatisation of quantum mechanics this question is particularly important because systems or mathematical structures we may usually associate to classical physics may have unexpected quantum properties. So, quantum mechanics has several defining principles, two of the most important being: its probabilistic nature: the outcomes generated by quantum mechanical calculations are always probabilistic, even if that probability gets very close to 0 or 1, and second, there is no pre-determined uniquely defined state of a system in the absence of an observable capable of asking a question about that property of the system. This property makes the probability amplitude and hence the wavefunction the maximal knowledge attainable about the system in a given context. Essentially, that means the quantitative properties of a state of a system are generally not uniquely defined, and sometimes are not even defined at all in the absence of an observable to be associated to a method of determining them. Moreover, not any observable is capable of fully determining the numerical value of a property of the state of a system, and moreover, some observables may not be compatible with respect to a given property. As a result, some of the matrix valued observables of quantum mechanics may not commute with each other. As a result, the wavefunction is, somehow unfortunately, called ”non-realistic”, and that is unfortunate because there is nothing ”unreal” about the wavefunction, in fact it provides the maximum of information about the system, and in many cases far more than a classical system could ever hope to achieve. However, a wavefunction ought to be interpreted as a catalogue of maximal knowledge that can be obtained about a property of the system to be analysed. As a consequence of this property, the rules of obtaining the probabilistic outcomes of quantum mechanics change. We therefore have to employ a method of obtaining a module of a complex number, from a fundamentally complex quantity, the wavefunction, that needs to be complex in order to account for correlations in our knowledge about the system. This in turn will generate the well known interference patterns in the probabilistic outcomes. That last property of quantum mechanics is Born’s rule.
@@
There exists a somehow widespread assumption that quantum phenomena must be either very small, or happen at low enough energy, or be small corrections of the order of $O(\hbar)$ to classical phenomena and that there is some form of ”gap” between a so called ”quantum world” and the ”real” or ”every-day” world. My thesis is that these assumptions are wrong. The approaches to quantum mechanics phenomena initiated up to now are based on controlling the interference of the ”catalogue of knowledge” wave patterns (basically, a sort of wavefunction engineering). Those methods are extremely susceptible to quantum decoherence, namely extended interference with the environment which leads to an exponential increase of the number of terms arising in the ”catalog of knowledge”, to the point where the sheer complexity of the problem makes the returns impossible to be understood in any meaningful sense.
@@
One direction of research nowadays is to find some classical underpinnings for quantum mechanics. This research is an important part of the so called ”foundational” research on quantum mechanics. While the common approach is to lift or extended the observables of a system from the uniquely defined classical real numbers to matrix valued observables that encode several possible outcomes without the possibility of predicting which one will occur from the data about the system itself, some quantum foundational research tries to find a classical (therefore well determined and unique) way of characterising the system based on some form of hidden classical variables. My approach is the exact opposite of this path. Instead of searching for some underpinning classical structure to quantum mechanics, my view is that even systems that appear as classical, may still have hidden quantum properties, due for example to the quantum nature of the transformation maps between the objects of the specific category considered. The same procedure that provides us with a quantum structure over the space of objects can provide us with a quantum structure over the space of maps between those objects in a categorical sense. If, for example, we employ geometric quantisation, following the prescriptions of pre-quantisation, polarisation or splitting, and half-form correction, we obtain (in most cases) a well defined structure on the objects (states). We could perform the same on the morphisms (or maps) of the category and obtain a quantum structure for those. In fact, it turns out that quantum information carried on the object space is found back also in the space of morphisms, and sometimes quantum properties of morphisms (like the cartesian non-separability leading to entanglement) may appear in the space of morphisms and have an effect on the space of objects as well.
@@
The assumption I wish to challenge is that quantum phenomena are strictly the result of objects being ”quantum” as opposed to ”classical”. Instead of a classical interaction between systems having well determined states, conventionally, one is expected to transition to a generalised interaction in which some classical object is replaced by a ”quantum” object in which not all of its properties are uniquely defined. This is the origin of integrating over inner loop momenta in Feynman diagrams or the construction of the Feynman path integral. This is of course correct, but it is not the only way quantum phenomena may emerge. Indeed, I wish to show that quantum phenomena can emerge strictly from the rules used to describe the dynamics of otherwise classical systems. Those rules themselves, seem more fundamental and can make the system generate quantum (or at least, almost quantum) behaviour. In this sense, even if the universe started with strictly classical objects and information (which is probably not the case), there exist certain rules of manipulating this classical information that will result in quantum behaviour. The quantum nature seems therefore embedded in the very rules of the system’s dynamics and does not need a separation between ”quantum” and ”classical” objects. The rules can be seen in categorical terms as maps, and the view drafted above would focus on the maps instead of the objects of a category.
@@
There are several macroscopic systems that are generally assumed to be classical, that by the way in which they work, have structures that give them some quantum properties, as far as the definitions above are concerned. One of these systems is the neural network. It is important to mention that there have been essays in which one speculated about the possibility of generating quantum computing processes in big enough biological neural networks by focusing on some detailed mechanism in some molecular substructure of the neural network [1]. I think such an approach is not feasible on a large scale (although one has discovered entanglement of spins in some animal brains [2,3] and substantial progress is being done in understanding macroscopic quantum phenomena from this point of view). Those methods are extremely susceptible to decoherence. However, one may ask what if the large structure of a network has properties that at least in part reflect the fundamental axioms of quantum mechanics? There have also been essays in constructing so called quantum neural networks [4]. Those networks have the links represented by quantum processes implemented by unitary transformations, turning the neural network into a large quantum circuit. Those methods are promising and it would be amazing if they will be achieved. However, there is still the problem of decoherence that will require such networks to be extremely robust, unfortunately a property still beyond today’s reach. It has however been noticed that classical neural networks do a great job at solving many body quantum problems with strong coupling [5], that are fundamentally quantum and expected to be properly solved on quantum computers. One role neural networks play in solving strongly correlated quantum systems is through simplifying the wavefunction by providing us with a more compact representation thereof. Another role played by neural networks is to improve the understanding of quantum dynamics [12-16]. What are the properties of a classical neural network that would make it suitable for solving such complex quantum problems? While it is true that classical neural networks seem to be very suitable for simulating quantum processes, I wish to ask in this article whether this is due to the fact that the rules of functioning of a neural network that appears classical from the perspective of its objects (namely neurons) has in fact hidden quantum properties at the level of the morphisms (maps) between those objects.
@@
Indeed, various quantum properties appear to be general properties not only of physical particles, but also of a series of mathematical structures that have a broad applicability. Non-separability, as encountered in the definition of entanglement, is also found, under the guise of non-cartesianity in the construction of fibre bundles, in which basically the bundle measures the global failure of maintaining the local cartesianity of a certain pairing. A non-trivial fibre bundle is defined to characterise precisely such a failure of the cartesian product between the fibre and the base to expand globally. This is just a mathematical example that could make the reasoning I try to follow here clearer: quantum properties are not bound to particles, but can also characterise various rules by which dynamical systems function. Such quantum properties make them function differently compared to their classical counterparts. Of course, this doesn’t mean a classical fibre bundle is quantum. We need additional structure for this. We need some type of quantum polarisation (a la geometric quantisation) and some form of half form correction, both general properties of quantisation, but there exists a deep similarity between the category of cobordisms whose objects are (n-1)-dimensional manifolds and its morphisms are n-dimensional cobordisms and the catgeory of Hilb whose objects are Hilbert spaces used to describe states and the morphisms are linear operators describing ”processes” [32]. With this relation, the cartesian non-separability property that leads to entanglement in quantum mechanics has a similar role in other situations that involve non-separability as is the case in fibre bundles.
Therefore a physical feature emerging from the quantum properties above is the inseparability of state spaces of quantum states [8]. This obstruction to cartesian pairing has been called entanglement, and it is advisable to keep that notion even when applied to more abstract mathematical objects [9], [32].
@@
In this article I show that a classical neural network can indeed be represented by a non-cartesian monodromic category and that there exists a functor that links it to the $Hilb$ category. By doing so, a normal classical neural network will show properties of quantum entanglement at a large scale in the same way in which other quantum systems show. However, being not quantum (i.e. highly decoherent) at the level of the constituent objects, it all depends on how we interpret the results of such a network. This means that the quantum nature emerges from the rules we impose on how the network should work, not from the actual quantum nature of its constituents, but finally, there should be no difference between a ”natural” quantum system made out of fundamental constituents that are ”quantum”, and one that is forced to function by rules (maps or morphisms of the category that are quantum) that result in an overall quantum behaviour of the network. The parameters of the network can be encoded in a Hilbert Kernel, where the respective parameters play the role of linear superpositions of the learned results, similar to the quantum superpositions allowed by the complex phases of a wavefunction. Therefore, there appears to be a strong analogy (if not more, a functorial connection) between entangled quantum states and neural networks and between the category of Hilbert spaces and the supposed category of neural networks.
@@
The distinction between a neural network as an ”unphysical” computational phenomenon as opposed to any other ”physical” phenomenon seems rather artificial. Any structure, mathematical or physical can be characterised as ”quantum” either because it has some intrinsic quantum nature (making some of the properties we expect it to have, undefined or not uniquely defined) or because the rules it must obey have to be such that the quantum nature emerges (becoming quantum due to the maps instead of objects). It is worth noting that recent research has succeeded in implementing a physical feed-forward neural network based on a series of physical phenomena that can be to some extent tuned to have only classical behaviour. Indeed it may be possible to experimentally test the morphisms origin of some quantum behaviour of a neural network by looking at a physical implementation of a neural network that appears to be classical but may gain quantum properties due to its inherent morphisms as such physical neural networks have been constructed [17]. Until then, I can offer a numerical simulation showing that indeed this is the case, namely that provided with proper input, a physical, classical neural network can entangle the particles that it receives as input.
@@
A classical neural network has certain properties that could be assimilated as quantum to some extent. While we would not propagate a quantum state through a series of quantum devices forming a hard to maintain network, we would use a set of classical neurons that, acting together, would be able to produce a state that could be classified as quantum for all practical purposes. Let us consider a neural network, made out of neurons firing according to classical rules. Let us also consider several layers of such neurons, as can be imagined in a deep learning network. The parameters of the neurons are bound together in a series of linear combinations in which the activation function plays the role of a non-linear contributor to the adaptability of the network. The relation among the neurons however is always linear. A signal is transmitted to the network as an input, which is passed then through the network and a gradient descent backpropagation method is used to train the linearly connected neurons forming the network to the desired structure. After defining a loss function, backpropagation is required to transmit the end-point error back to the input node and to optimise the parameters in the process. This process allows the network to access, via its non-linear activation function and particularly through its loss function, the non-local structure of the features it explores. This is also why neural networks are so good at classifying non-local classification problems and are finally used as approximations for highly entangled quantum states. But are classical neural networks just approximations of quantum systems, or are they, in some sense, quantum themselves, not through the interactions that make them function, but through their global structure and their backpropagation and loss function features? Usually, in a neural network, the optimised result is not easily associated with the input, which is why the artificial neural networks have the well known problem of inferring causal connections. This may not be surprising given that the end-result is obtained through a process of extensive linear superposition of learned data and then optimised via a loss function that introduces the learned data to the global structure of the problem it is asked to solve. This is what an entangling gate would do. A Hadamard gate would create a linear superposition of the initial state. This is what occurs to our initial data via the network linear superposition. A CNOT gate would combine two states, switching the result of one when the other is in a specified state. This is generally the result of our loss functions. The only difference is that we are not a-priori acting on quantum states but on suitably represented initial learning data. It all depends on what meaning we give to that data and how we interpret the output then, to make it a quantum analogue or a simple classical superposition.
@@
$K(Z,t)=H(z,t)+\frac{dF(q,P,t)}{dt}=H(\Psi(Z,t),t)+\frac{dF(\Psi_{1}(Z,t),P,t)}{dt}$ (52)

### v4

@@ Abstract
Neural networks are being used to improve the probing of the state spaces of many particle systems as approximations to wavefunctions and in order to avoid the recurring sign problem of quantum monte-carlo. One may ask whether the usual classical neural networks have some actual hidden quantum properties that make them such suitable tools for a highly coupled quantum problem. I discuss here what makes a system quantum and to what extent we can interpret a neural network as having quantum remnants. I suggest that a system can be quantum both due to its fundamental quantum constituents and due to its dynamics, therefore, we can obtain entanglement both due to the quantum constituents’ nature and due to the functioning rules, or, in category theory terms, both due to the quantum nature of the objects of a category and of the maps. From a practical point of view, I suggest a possible experiment that could extract entanglement from the dynamics (maps) of an otherwise cvasi-classical (from the point of view of the constituents) neural network.
@@ I Introduction
Neural networks are nowadays regarded as computational tools that have a broad set of applications from quantum chemistry to solving logistics problems. However, in all biological circumstances, neural networks appear as special materials that have a flexible design reaching down to the cellular level and with an impact on various other tissues like muscles, vascular systems, etc. As such, a neural network is a special material with surprising physical, chemical, and, as I wish to present in this article, quantum properties.
@@
The goal of this article is to show that the dynamics of a cvasi-classical neural network may have underlying quantum properties that can be amplified by its dynamics. It must be noted that the neural network referred to in this article is not a numerical simulation, as a simulation would ultimately, by design, have no complex phase and hence would not present quantum effects. My discussion is based on a real physically implemented cvasi-classical neural network. The main observation is that such a neural network may generate quantum effects due to the dynamics (or, in a categorical sense, due to the maps between the objects, not the objects themselves). The end result is that a cvasi-classical neural network, due to its dynamics, could generate entangled states and could be used for entanglement generation and extraction. As such, this is of particular interest to subjects like quantum biology or to describing quantum phenomena in natural neural networks where the usual criteria of temperature and scale would otherwise block such effects. Probably the best way of starting this article is to ask a question: what does it mean to be quantum? Lacking a proper axiomatisation of quantum mechanics this question is particularly important because systems or mathematical structures we may usually associate to classical physics may have unexpected quantum properties. From another point of view, systems usually modelled by means of strictly classical rules, may, upon physical implementation present quantum behaviour that escaped our previous analysis. Quantum mechanics has several defining principles, two of the most important being: its probabilistic nature: the outcomes generated by quantum mechanical calculations are always probabilistic, even if that probability gets very close to 0 or 1, and second, there is no pre-determined uniquely defined state of a system in the absence of an observable capable of asking a question about that property of the system. This property makes the probability amplitude and hence the wavefunction of the system represent the maximal knowledge attainable about the system in a given context. Essentially, that means the quantitative properties of a state of a system are generally not uniquely defined, and sometimes are not even defined at all in the absence of an observable to be associated to a method of determining them. Moreover, not any observable is capable of fully determining the numerical value of a property of the state of a system, and some observables may not be compatible with respect to a given property. As a result, some of the matrix valued observables of quantum mechanics may not commute with each other. Due to these properties, the wavefunction is, somehow unfortunately, called ”non-realistic”, and that is unfortunate because there is nothing ”unreal” about the wavefunction, in fact it provides the maximum of information about the system, and in many cases far more than a classical system could ever hope to achieve. A wavefunction ought to be interpreted as a catalogue of maximal knowledge that can be obtained about a property of the system to be analysed. As a consequence, the rules of obtaining the probabilistic outcomes of quantum mechanics change. We therefore have to employ a method of obtaining a module of a complex number, from a fundamentally complex quantity, the wavefunction. This wavefunction needs to be complex in order to account for correlations in our knowledge about the system. In turn, by this method we will generate the well known interference patterns in the probabilistic outcomes of quantum mechanics. That last property of quantum mechanics is Born’s rule.
@@
There exists a somehow widespread assumption that quantum phenomena must be either very small, or happen at low enough energy, or be small corrections of the order of $O(\hbar)$ to classical phenomena and that there is some form of ”gap” between a so called ”quantum world” and the ”real” or ”every-day” world. My thesis is that these assumptions are wrong. The approaches to quantum mechanics phenomena initiated up to now are based on controlling the interference of the ”catalogue of knowledge” wave patterns (basically, a sort of wavefunction engineering). Those methods are extremely susceptible to quantum decoherence, namely extended interaction and entanglement with the environment which leads to an exponential increase of the number of terms arising in the ”catalog of knowledge” describing the environment and the system. If one approaches quantum phenomena from a categorical perspective, one notices that the maps between quantum objects may as well have quantum properties, may become non-cartesian, and can produce and exploit entanglement. The maps between quantum objects are associated to the specific dynamics through which the system goes, and this very dynamics can produce, enhance, and utilise entanglement.
@@
One direction of research nowadays is to find some classical underpinnings for quantum mechanics. This research is an important part of the so called ”foundational” research on quantum mechanics. While the common approach is to lift or extend the observables of a system from the uniquely defined classical real numbers to matrix valued observables that encode several possible outcomes without the possibility of predicting which one will occur from the data about the system itself, some quantum foundational research tries to find a classical (therefore well determined and unique) way of characterising the system based on some form of hidden classical variables. My approach is the exact opposite of this path. Instead of searching for some underpinning classical structure to quantum mechanics, my view is that even systems that appear as classical, may still have hidden quantum properties, due, for example to the quantum nature of the transformation maps between the objects of the specific category considered. In particular, systems assumed to be classical, like a neural network, despite being simulated as classical objects in conventional computers, when physically realised, may exhibit quantum phenomena as cvasi-classical systems. In particular, the non-linear dynamics of neural networks may in fact amplify quantum behaviour.
The same procedure that provides us with a quantum structure over the space of objects can provide us with a quantum structure over the space of maps between those objects in a categorical sense. If, for example, we employ geometric quantisation, following the prescriptions of pre-quantisation, polarisation or splitting, and half-form correction, we obtain (in most cases) a well defined structure on the objects (states). We could perform the same on the morphisms (or maps) of the category and obtain a quantum structure for those. In fact, it turns out that quantum information carried on the object space is found also in the space of morphisms, and sometimes quantum properties of morphisms (like the cartesian non-separability leading to entanglement) may appear in the space of morphisms and have an effect on the space of objects as well.
@@
The assumption I wish to challenge is that quantum phenomena are strictly the result of objects being ”quantum” as opposed to ”classical”. Instead of a classical interaction between systems having well determined states, conventionally, one is expected to transition to a generalised interaction in which some classical object is replaced by a ”quantum” object for which not all of its properties are uniquely defined. This is the origin of the requirement of integrating over inner loop momenta in Feynman diagrams or the construction of the Feynman path integral. This is of course correct, but it is not the only way quantum phenomena may emerge. Indeed, I wish to show that quantum phenomena can emerge strictly from the rules used to describe the dynamics of cvasi-classical systems. Those rules themselves, seem more fundamental and can make the system generate quantum (or at least, almost quantum) behaviour. In this sense, even if the universe started with strictly classical objects and information (which is probably not the case), there exist certain rules of manipulating this classical information that will result in quantum behaviour. The quantum nature seems therefore embedded in the very rules of the system’s dynamics and does not need a separation between ”quantum” and ”classical” objects. The rules can be seen in categorical terms as maps, and the view drafted above would focus on the maps instead of the objects of a category. In the case of a cvasi-classical neural network, those rules would be those of the dynamical behaviour of the network.
@@
There are several macroscopic systems that are generally assumed to be classical, that by the way in which they work, have structures that give them some quantum properties, as far as the definitions above are concerned. One of these systems is the neural network. It is important to mention that there have been essays in which one speculated about the possibility of generating quantum computing processes in big enough biological neural networks by focusing on some detailed mechanism in some molecular substructure of their neural network [1]. I think such an approach is not feasible on a large scale (although one has discovered entanglement of spins in some animal brains [2,3] and substantial progress is being done in understanding macroscopic quantum phenomena from this point of view). Those methods are extremely susceptible to decoherence. However, one may ask what if the large structure of a network has properties that at least in part reflect the fundamental axioms of quantum mechanics? There have also been essays in constructing so called quantum neural networks [4]. Those networks have the links represented by quantum processes implemented by unitary transformations, turning the neural network into a large quantum circuit. Those methods are promising and it would be amazing if they will be achieved. However, there is still the problem of decoherence that will require such networks to be extremely robust, unfortunately a property still beyond today’s reach. It has however been noticed that classical neural networks do a great job at solving many body quantum problems with strong coupling [5], that are fundamentally quantum and expected to be properly solved on quantum computers. One role neural networks play in solving strongly correlated quantum systems is through simplifying the wavefunction by providing us with a more compact representation thereof. Another role played by neural networks is to improve the understanding of quantum dynamics [12-16]. What are the properties of a classical neural network that would make it suitable for solving such complex quantum problems? While it is true that classical neural networks seem to be very suitable for simulating quantum processes, I wish to ask in this article whether this is due to the fact that the rules of functioning of a neural network that appears classical from the perspective of its objects (namely neurons) has in fact hidden quantum properties at the level of the morphisms (maps) between those objects.
@@
Indeed, various quantum properties appear to be general properties not only of physical particles, but also of a series of mathematical structures that have a broad applicability. Non-separability, as encountered in the definition of entanglement, is also found, under the guise of non-cartesianity in the construction of fibre bundles, in which basically the bundle measures the global failure of maintaining the local cartesianity of a certain pairing. A non-trivial fibre bundle is defined to characterise precisely such a failure of the cartesian product between the fibre and the base to expand globally. This is just a mathematical example that could make the reasoning I try to follow here clearer: quantum properties are not bound to particles, but can also characterise various rules by which dynamical systems function. Such quantum properties make them function differently compared to their classical counterparts. Of course, this doesn’t mean a classical fibre bundle is quantum. We need additional structure for this. We need some type of quantum polarisation (a la geometric quantisation) and some form of half form correction, both general properties of quantisation, but there exists a deep similarity between the category of cobordisms whose objects are (n-1)-dimensional manifolds and its morphisms are n-dimensional cobordisms, and the catgeory of Hilb whose objects are Hilbert spaces used to describe states and the morphisms are linear operators describing ”processes” [32]. With this relation, the cartesian non-separability property that leads to entanglement in quantum mechanics has a similar role in other situations that involve non-separability, as is the case in fibre bundles. From a practical perspective, it is important to know that such entanglement exists in a category related by a functor to the initial one, as it would be possible to harness it as a quantum resource and use it in circumstances that are apparently remote from the initial categorical structure. In particular, an apparently classical neural network may extract its entanglement from an associated category and behave like a quantum system when its dynamics is analysed. A wavefunction, or a probability amplitude (a precursor to the probability we are used to in classical physics) can be associated not only to the objects of the category, but also to the maps between those objects, and hence superposition and even entanglement described at the level of the wavefunction associated to the quantum maps may prove to be a new quantum resource that could be harnessed by physical neural networks. This would lead to the possibility of having quantum computation like effects occurring even in the most noisy and hot environments, like the ones existing in animal brains. A physical feature emerging from the quantum properties above is the inseparability of state spaces of quantum states [8]. This obstruction to cartesian pairing has been traditionally called entanglement, and it is advisable to keep that notion even when applied to more abstract mathematical objects [9], [32].
@@
In this article I show that a (cvasi)-classical neural network can indeed be represented by a non-cartesian monodromic category and that there exists a functor that links it to the $Hilb$ category. By doing so, a normal (cvasi)-classical neural network will show properties of quantum entanglement at a large scale in the same way in which other quantum systems show. However, being highly decoherent at the level of the constituent objects, it will be up to the maps to act as a resource for quantum entanglement. This means that the quantum nature emerges from the rules we impose on how the network should work, not from the actual quantum nature of its constituents, but finally, there should be no difference between a ”natural” quantum system made out of fundamental constituents that are ”quantum”, and one that is forced to function by rules (maps or morphisms of the category that are quantum) that result in an overall quantum behaviour of the network. The parameters of the network can be encoded in a Hilbert Kernel, where the respective parameters play the role of linear superpositions of the learned results, similar to the quantum superpositions allowed by the complex phases of a wavefunction. Therefore, there appears to be a strong analogy (if not more, a functorial connection) between entangled quantum states and neural networks and between the category of Hilbert spaces and the supposed category of neural networks.
@@
The distinction between a neural network as an ”unphysical” computational phenomenon as opposed to any other ”physical” phenomenon seems rather artificial. Any structure, mathematical or physical can be characterised as ”quantum” either because it has some intrinsic quantum nature (making some of the properties we expect it to have, undefined or not uniquely defined) or because the rules it must obey have to be such that the quantum nature emerges (becoming quantum due to the maps instead of objects). It is worth noting that recent research has succeeded in implementing a physical feed-forward neural network based on a series of physical objects that can be to some extent tuned to have only classical behaviour. Indeed it may be possible to experimentally test the morphisms origin of some quantum behaviour of a neural network by looking at a physical implementation of a neural network that appears to be classical from the perspective of the objects involved, but may gain quantum properties due to its inherent morphisms. Such physical neural networks have been constructed [17]. Until such an experiment can be performed, I can offer a numerical simulation showing that indeed this is the case, namely that provided with proper input, a physical, classical neural network can entangle the particles that it receives as input.
@@
A (cvasi)-classical neural network has certain properties that could be assimilated as quantum to some extent. While we would not propagate a quantum state through a series of quantum devices forming a hard to maintain network, we would instead use a set of classical neurons that, acting together, would be able to produce a state that could be classified as quantum for all practical purposes. Let us consider a neural network, made out of neurons firing according to classical rules. Let us also consider several layers of such neurons, as can be imagined in a deep learning network. The parameters of the neurons are bound together in a series of linear combinations in which the activation function plays the role of a non-linear contributor to the adaptability of the network. The relation among the neurons however is always linear. A signal is transmitted to the network as an input, which is passed then through the network and a gradient descent backpropagation method is used to train the linearly connected neurons forming the network to the desired structure. After defining a loss function, backpropagation is required to transmit the end-point error back to the input node and to optimise the parameters in the process. This process allows the network to access, via its non-linear activation function and particularly through its loss function, the non-local structure of the features it explores. This is also why neural networks are so good at classifying non-local classification problems and are finally used as approximations for highly entangled quantum states. But are classical neural networks just approximations of quantum systems, or are they, in some sense, quantum themselves, not through the interactions that make them function, but through their global structure and their backpropagation and loss function features (mainly their dynamics maps)? Usually, in a neural network, the optimised result is not easily associated with the input, which is why the artificial neural networks have the well known problem of inferring causal connections. This may not be surprising given that the end-result is obtained through a process of extensive linear superposition of learned data and then optimised via a loss function that introduces the learned data to the global structure of the problem it is asked to solve. This is what an entangling gate would do. A Hadamard gate would create a linear superposition of the initial state. This is what occurs to our initial data via the network linear superposition. A CNOT gate would combine two states, switching the result of one when the other is in a specified state. This is generally the result of our loss functions. The only difference is that we are not a-priori acting on quantum states but on suitably represented initial learning data. It all depends on what meaning we give to that data and how we interpret the output then, to make it a quantum analogue or a simple classical superposition.
@@
$\begin{array}[]{c}K(Z,t)=H(z,t)+\frac{dF(q,P,t)}{dt}=H(\Psi(Z,t),t)+\\ \\ +\frac{dF(\Psi_{1}(Z,t),P,t)}{dt}\\ \end{array}$ (52)